OpenStudy (anonymous):

Find the product of (x + 3)^2. Is this the answer? x^2 + 6x + 9

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

(x + 3)^2 = x(x+3)+3(x+3) = x^2 + 3x + 3x + 9 = x^2 + 6x + 9

OpenStudy (anonymous):

Okay so could you help me with this? Find the product of (x + 3)(x - 3).

OpenStudy (anonymous):

\[(a+b)^2 = a^2 + 2ab + b^2 \\ (a-b)^2 = a^2 -2ab + b^2 \\ a^2-b^2 = (a+b)(a-b)\]